OpenStudy (k8lyn911):

Calculus II Online Homework (**Due at midnight**): Find the absolute maximum and absolute minimum values of f on the given interval. f(x)= x^(2) e^(-x/2) [-1,8]

OpenStudy (rizags):

max{f(x) = e^(-x/2) x^2} = 16/e^2 at x = 4

OpenStudy (rizags):

min{f(x) = e^(-x/2) x^2} = 0 at x = 0

OpenStudy (rizags):

not sure about these though, been a while since calculus

OpenStudy (rizags):

anyone can check?

OpenStudy (k8lyn911):

I typed them in, and they're right. :) Thank you so much!

OpenStudy (rizags):

fan me!

OpenStudy (k8lyn911):

Will do! :)

OpenStudy (anonymous):

That is correct